OpenStudy (anonymous):

The conical hat shown here is a traditional sun-protective straw hat originating in eastern Asia. What is the approximate outer surface area of one of these hats that has a radius of 9 inches and a slant height of inches? Keep in mind that there is an opening in the hat where the base surface would normally be. Show the steps of your solution and explain your work.

OpenStudy (anonymous):

Slant height is 12 in *****

OpenStudy (anonymous):

@AnimalAin

OpenStudy (anonymous):

how many inches is the slant height?

OpenStudy (anonymous):

do you know the formula for surface area of a cone?

OpenStudy (anonymous):

sa cone : pie x r x s + pie x radius squared.

OpenStudy (anonymous):

slant height is 12 inches .

OpenStudy (anonymous):



OpenStudy (anonymous):

you first need to find the height of the cone using the Pythagorean Theorem

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\] a=9, b=height, c=12

OpenStudy (anonymous):

solve for b

OpenStudy (anonymous):

7.9 ?

OpenStudy (anonymous):

7.94 just to be safe-oops you didn't need to find the height haha that's for the volume..sorry! Now plug everything into the SA formula \[\pi(9)(12)+\pi(9)\]

OpenStudy (anonymous):

\[\pi(9)^{2}\]

OpenStudy (anonymous):

First find it in terms of pi, it will be easier

OpenStudy (anonymous):

so i didnt need to do the pythagoream theorm ?

OpenStudy (anonymous):

sa = 367.5

OpenStudy (anonymous):

leave it as \[189\pi\] for now we have to subtract the area of the base now

OpenStudy (anonymous):

\[189\pi-\pi(9)^{2}\]

OpenStudy (anonymous):

108 pie ?

OpenStudy (anonymous):

@J.L. ??

OpenStudy (ash2326):

Yeah, you're right

OpenStudy (ash2326):

since a hat doesn't have a circular base, it's open

OpenStudy (anonymous):

so its 367.5 for the outer surface area?

OpenStudy (anonymous):

Your answer should be 339.292

OpenStudy (anonymous):

Consider it approximated by a triangle that is 12 tall and has a base of 18 pi. That might not get you a ten-digit correct answer, but will give a satisfactory approximation.

OpenStudy (anonymous):

By the way, if you say the cut and flattened hat forms a sector of a 12 inch radius circle that is three fourths of the circle. The area of that sector is 108 pi, exactly the same as my approximation above.

OpenStudy (anonymous):

???